OpenStudy (anonymous):

Solve: 2(x + 2) − (x + 1) = 8

OpenStudy (anonymous):

2x+4-x-1=8 2x-x+4-1-8=0 now simplify it..

OpenStudy (anonymous):

so: \[x+(-5) ?\]

OpenStudy (anonymous):

yes ryt..

OpenStudy (anonymous):

@vetdoc you can't not abandon the right side .It's a Equation

OpenStudy (anonymous):

Dont I do the distributive property first?

OpenStudy (anonymous):

Yeah, you have to take in consideration the other side, but you are right x - 5 =8 Now, can you find x?

OpenStudy (anonymous):

so then I would do: \[x=13?\]

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

Oh okay! Thanks I got it now!

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

@arch1 no, it simplifies to x-5 = 8 so you can add the 5 to both sides to isolate x, in which you get x = 5+18, or x=13

OpenStudy (anonymous):

@thechocoluver445 :hw wil u solve this 2x+4-x-1=8?

OpenStudy (anonymous):

 O, I c what you meant now. I was going of the x-5=8 thing above, my bad :P